OpenStudy (anonymous):

Identify which equation is a quadratic equation. y = −6x − 3 y + x2 = (x − 1)(x + 7) y − 9 = (x − 1)(x2 + 4) y = 5x(x + 3) − 8

OpenStudy (anonymous):

y = 5x(x + 3) − 8 because y= 5x^2 + 15x about the second one y+x^2 = x^2+6x-7, x power 2 would be omitted!

OpenStudy (anonymous):

Thanksss so much . c: